OpenStudy (anonymous):

The perimeter of a rectangle is 96 ft. The ratio of its length to its width is 7 : 5. What are the dimensions of the rectangle? A. 30 ft by 18 ft B. 28 ft by 20 ft C. 24 ft by 24 ft D. 22 ft by 26 ft

OpenStudy (anonymous):

we have : W+L=96/2 K ? And have a Proportion too ! So have : W+L=48 7+1=8 have : w_____________1_____________x L-----------7-------------y perimeter____8________48 we have 8*6=18 so have : 7*6=42 1*6=6 Got it ?:)

OpenStudy (anonymous):

thanks! :">

OpenStudy (anonymous):

welcome :)